I also need help with theese three if you dont mind me asking for it.
4(3 − 2) = 4(3) − 4(2)
9 + 8 = 8 + 9
4 + (2 + 6) = (4 + 2) + 6

Answer by KnightOwlTutor(293)

4(3 − 2) = 4(3) − 4(2) Distributive-You can muliply through the difference
9 + 8 = 8 + 9 =Communitive -the order doesn't matter
4 + (2 + 6) = (4 + 2) + 6=Associative -the grouping of a sum does not alter its value